OpenStudy (freemap):

The table shows the types of popcorn buckets sold at a concession stand for one day. Butter No Butter Small 182 111 Medium 140 26 Large 97 44 What percent of the buttered popcorn buckets were large? Round to the nearest percent.

OpenStudy (freemap):

Butter No Butter Small 182 111 Medium 140 26 Large 97 44

OpenStudy (michele_laino):

buttered popcorns are \(182+140+97=419\)

OpenStudy (freemap):

There's supposed to be a line dividing butter and no butter okay just got your post yes

OpenStudy (michele_laino):

so the requested percentage, is: \[\huge \% = \frac{{97}}{{419}} \times 100 = ...?\]

OpenStudy (freemap):

23.150

OpenStudy (michele_laino):

yes!

OpenStudy (freemap):

Round to the nearest percent?

OpenStudy (michele_laino):

it is \(23\%\)

OpenStudy (freemap):

Thank you

OpenStudy (michele_laino):

:)